OpenStudy (anonymous):

product of all factors of number N is equal to tenth power of the number . then N cant be divisible by 90 , 200 , 126 or 210?

OpenStudy (shubhamsrg):

We need to prove it ?

OpenStudy (anonymous):

i give 4 numbers, you have to choose the correct option. basically the question is: product of all factors of number N is equal to tenth power of the number . then N cant be divisible by a. 90 b. 200 c. 126 d. 210

OpenStudy (shubhamsrg):

Is the ans 210 ?

OpenStudy (shubhamsrg):

By my work, my best bet will be 210

OpenStudy (anonymous):

well i dont have the solutions right now but i too got 210 by looking at its factor. how did you get it?

OpenStudy (shubhamsrg):

Let N be non perfect square. Let there be m factors of N, then product of all factors of N = N^(m/2) We are given N^(m/2) = N^10 hence m=20 factoring 20 -> (1*20) = (2*10) = (4*5) = (2*2*5) There can thus be max 3 prime factors of N Only 210 has 4 prime factors, rest all have 3 . And hence the asn

OpenStudy (perl):

why is N= N^(m/2) and how do you know N is a non perfect square

OpenStudy (perl):

im going to bed i will see you tomorrow

OpenStudy (shubhamsrg):

It is n^10 = n^(m/2)

OpenStudy (perl):

where does m come from?

OpenStudy (perl):

please explain more , thanks

OpenStudy (shubhamsrg):

And Its and assumption, you'll have to take another case where N is a perfect sq.

OpenStudy (shubhamsrg):

an*

OpenStudy (shubhamsrg):

Lets take an example, say 20

OpenStudy (shubhamsrg):

factors = 1,2,4,5,10,20

OpenStudy (shubhamsrg):

product = (1*20) * (2*10) * (4*5)

OpenStudy (perl):

there are more, 8,

OpenStudy (perl):

oh woops

terenzreignz (terenzreignz):

60 is such a number... as long as factors don't include the number itself.

terenzreignz (terenzreignz):

oh woops

OpenStudy (shubhamsrg):

so product = 20^3 = 20^(6/2)

OpenStudy (shubhamsrg):

Whats with "woops" today? o.O

terenzreignz (terenzreignz):

Too impulsive... working on a solution :P

OpenStudy (shubhamsrg):

So N^(m/2) is clear ?

terenzreignz (terenzreignz):

Trying a different approach... hang on.

OpenStudy (perl):

this is true for all numbers?

OpenStudy (shubhamsrg):

Which are not squares.

OpenStudy (shubhamsrg):

Alternatively, all sq. no. have odd no. of factors, otherwise, all other nos. have even no. of factors.

OpenStudy (shubhamsrg):

So 2 of those factors will always multiply together to give N. No. of pairs = m/2 Hence product = N^(m/2)

OpenStudy (perl):

thats not a rigorous argument, but ok, i guess this just popped up

OpenStudy (shubhamsrg):

What do you mean by not a rigorous argument ?

OpenStudy (perl):

doing one example doesnt mean its true for all non perfect squares

OpenStudy (shubhamsrg):

Lets try to generalize.

OpenStudy (shubhamsrg):

Let factors of N be p1, p2,p3 .. pm (i.e. m factors)

OpenStudy (shubhamsrg):

Where p1<p2<p3..<pm

OpenStudy (shubhamsrg):

p1 * pm = N

terenzreignz (terenzreignz):

None of which are 1 or N, right? You can be specific and let there be 2k factors, but

OpenStudy (shubhamsrg):

p2 * pm-1 = N

OpenStudy (shubhamsrg):

Nopes. 1 and N included.

OpenStudy (shubhamsrg):

If we exclude both, product will be N^(m-2)/2

terenzreignz (terenzreignz):

well, fine, but you can be specific and make it so that they are p1, p2,p3... pk, And then the rest are N/p1,N/p2,...N/pk Like that

OpenStudy (perl):

ok i think I see what you did

OpenStudy (shubhamsrg):

Glad you do.

OpenStudy (perl):

(2^0 * 5^0) * (2^2 * 5^1 (2^1 *5^0) * ( 2^1 * 5^1) ...

OpenStudy (shubhamsrg):

?

OpenStudy (perl):

how do you find N

OpenStudy (shubhamsrg):

We don;t have to find N, where did I find N

OpenStudy (perl):

you wrote There can thus be max 3 prime factors of N Only 210 has 4 prime factors, rest all have 3 .

OpenStudy (perl):

how did you reason that there can be max 3 prime factors

OpenStudy (shubhamsrg):

Yes. m =20 , i.e. N has total 20 no. of factors.

OpenStudy (shubhamsrg):

20 can be written as 1*20 2*10 4*5

OpenStudy (shubhamsrg):

or 2*2*5

OpenStudy (shubhamsrg):

Before this, you should be knowing that , say n = x^p . y^q . z^r , where x,y,z are prime nos, then total factors of n are (p+1)(q+1)(r+1)

OpenStudy (perl):

im not sure what they are asking. you didnt find N

OpenStudy (perl):

terenz, what are they asking?

terenzreignz (terenzreignz):

Something tells me I should stay out of this, but it's only asking which of those four numbers cannot divide N.

OpenStudy (shubhamsrg):

We are not supposed to find N, we are supposed to find which of the following (90,126,210,200) is not a factor of N

OpenStudy (perl):

ok

OpenStudy (shubhamsrg):

Are you following till now ?

OpenStudy (shubhamsrg):

Do you know about no. of factors thing ? Like I just said , (p+1)(q+1)(r+1) ?

OpenStudy (perl):

now sure how you got that

OpenStudy (shubhamsrg):

hmmm..

OpenStudy (perl):

nevermind, i see that

OpenStudy (perl):

thats just using number of choices

OpenStudy (shubhamsrg):

you sure?

OpenStudy (shubhamsrg):

Yep.

OpenStudy (perl):

yeah, im lost on the other stuff

OpenStudy (perl):

so we deduced that N has 20 factors, because N=N^(m/2) = N^(10) ---> m /2 = 10 , m = 20

OpenStudy (shubhamsrg):

yes.

OpenStudy (shubhamsrg):

20 must be of the form of (p+1)(q+1)(r+1)(s+1) ...

OpenStudy (shubhamsrg):

Where p,q,r,s.. are all natural numbers.

OpenStudy (perl):

N = x^p . y^q . z^r *... 20 = (p+1)(q+1)(r+1) ... ok

OpenStudy (shubhamsrg):

We also know 20 = 1*20 = 2*10 = 4*5 = 2*2*5 Not any more combos possible right ?

OpenStudy (perl):

right

OpenStudy (shubhamsrg):

Lets sat p<=q<=r<=s... On comparison, p can be 0 and then q = 19 (for 1*20) p can be 1, and q = 9 ( for 2*10) p can be 3 ,q=4 , (for 4*5) p can be 1, q=1 and r =2 (for 2*2*5)

terenzreignz (terenzreignz):

Heh... I give up @shubhamsrg Your way was probably the only way that's not beyond my own capabilities... There's no need to reinvent the wheel...

OpenStudy (shubhamsrg):

hmmm. I should take that as a compliment! :P

OpenStudy (perl):

r = 2 ?

OpenStudy (shubhamsrg):

sorry 4*

OpenStudy (perl):

ok, now how does that help to get the conclusion

terenzreignz (terenzreignz):

I literally made guesses on possible combinations of the first three prime numbers (my rationale was that 3 is the greatest natural number where 3! < 10) And figured out (experimentally, not logically) that (p+1)(q+1)(r+1) thing @shubhamsrg had already come up with earlier. I guess the least possible value for N is 240.

OpenStudy (perl):

let me think about this and come back tomorrow you also have to look at the perfect square case?

terenzreignz (terenzreignz):

Well, anyway, @shubhamsrg way of doing it had no flaw to it that I could spot, but apparently just needed more explaining :P

OpenStudy (shubhamsrg):

Now we know N = x^p . y^q . z^r there is no possibility of a "s"

OpenStudy (shubhamsrg):

Hence N can only we represented in terms of 3 prime factors. only 210 is the one which can not be represented by 3 prime numbers

terenzreignz (terenzreignz):

@perl Cannot be a perfect square if you accept that N=(x^p)(y^q)(z^r) As that would imply that p,q, and r are all even, implying (p+1)(q+1)(r+1) = 20 to be odd, a contradiction.

OpenStudy (shubhamsrg):

If we assume N is perfect sq, then product of factors will be N^((m+1)/2) => m=19 19 = 1*19 only hence N = x^18 (only) Hope you can make out where this will go.